# Each of those eight coalition logics is also determined by six other kinds of models

### v2

@@ 4.4 Representation of single-coalition-first action models by single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as follows:
@@
Let $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
@@
Let $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. It suffices to show $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\text{ and }\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\}$.
@@
Assume $\mathtt{suc}(s)=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\emptyset$ and $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$ and $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
@@
Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
@@
Let $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $\mathrm{X}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{1}}(s,\sigma_{a_{n}})$ and $\mathrm{X}\neq\emptyset$. Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\neq\emptyset$, …, $\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\neq\emptyset$. Let $i$ be such that $1\leq i\leq n$ and $\sigma_{a_{i}}=\alpha_{x-u-\mathrm{X}_{i}}$. By the definition of $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})$, $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})=\mathrm{X}_{i}$, $x=a_{i}$, and $u=s$. Then $\sigma_{a_{i}}=\alpha_{a_{i}-s-\mathrm{X}_{i}}$ and $\mathrm{X}_{i}\in\mathtt{nei}_{a_{i}}(s)$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$. Then, $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$.
@@
Let $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s,\mathrm{C})$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$ for some $\mathrm{X}_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{n}\in\mathtt{nei}_{a_{n}}(s)$. Then, $\alpha_{a_{1}-s-\mathrm{X}_{1}}\dots,\alpha_{a_{n}-s-\mathrm{X}_{n}}\in\mathtt{AC}$. Note $\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})=\mathrm{X}_{1}$, $\dots$, $\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{n}$. Let $\sigma_{\mathrm{C}}=\alpha_{a_{1}-s-\mathrm{X}_{1}}\cup\dots\cup\alpha_{a_{n}-s-\mathrm{X}_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}=\mathrm{X}$. Note $\mathrm{X}$ is not empty. Then, $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$. Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then, for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
@@ 5.1 Clear grand-coalition-first action models
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model.
We say $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
The following result offers two alternative definitions of clear grand-coalition-first action models.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
@@
For any set $\mathrm{X}$ of states, we say $\Delta\subseteq\mathcal{P}(\mathrm{X})$ is a general partition of $\mathrm{X}$ if (1) $\bigcup\Delta=\mathrm{X}$ and (2) for every $\mathrm{Y},\mathrm{Y}^{\prime}\in\Delta$, if $\mathrm{Y}\neq\mathrm{Y}^{\prime}$, then $\mathrm{Y}\cap\mathrm{Y}^{\prime}=\emptyset$. Note a general partition $\Delta$ is a partition if $\emptyset\notin\Delta$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the sucessor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
Then, if $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, then for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $t\in\mathtt{ST}$. Note $\mathtt{suc}(s)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$. It is easy to show $\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}=\mathtt{suc}(s)$. Let $\mathrm{X}_{1},\mathrm{X}_{2}\in\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ such that $\mathrm{X}_{1}\neq\mathrm{X}_{2}$. Then, $\mathrm{X}_{1}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{1})$ for some $\sigma_{\mathrm{C}}^{1}\in\mathtt{JA}_{\mathrm{C}}$ and $\mathrm{X}_{2}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{2})$ for some $\sigma_{\mathrm{C}}^{2}\in\mathtt{JA}_{\mathrm{C}}$. Note $\sigma^{1}_{\mathrm{C}}\neq\sigma^{2}_{\mathrm{C}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$, that is, $\mathrm{X}_{1}\cap\mathrm{X}_{2}=\emptyset$.
Note the other direction of the statement of this fact might not hold. Suppose $\mathtt{AG}=\{a\}$. Consider a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ such that:
@@
It is easy to check for all $\mathrm{C}\subseteq\mathtt{AG}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathtt{suc}(s)$. However, $\mathtt{G}\textbf{-}\mathtt{AM}$ is not clear.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a clear grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $\mathtt{suc}$ be the successor function of $\mathtt{G}\textbf{-}\mathtt{AM}$.
@@
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.

### v3

@@ 4.4 Representation of single-coalition-first action models by single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ as follows:
@@
    • for every $s\in\mathtt{ST}$, $\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\mathtt{nei}_{\emptyset}(s)$.
@@
    • for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{\mathrm{C}}}|_{a})\mid a\in\mathrm{C}\}$.
It is easy to check that for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$. Then $\mathtt{S}\textbf{-}\mathtt{AM}$ is a single-coalition-first action model.
@@
Let $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. It suffices to show $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\text{ and }\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\neq\emptyset\}$.
@@
Assume $\bigcup\mathtt{nei}_{\emptyset}(s)=\emptyset$. Then, $\mathsf{out}_{\emptyset}(s,\emptyset)=\emptyset$. Note for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, for all $a\in\mathtt{AG}$, $\sigma_{a}\in\mathtt{JA}_{a}$ and $s\in\mathtt{ST}$, $\mathsf{out}_{a}(s,\sigma_{a})=\emptyset$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\emptyset=\mathtt{nei}_{\mathrm{C}}(s)$.
Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Note $\mathtt{AE}_{\mathrm{C}}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$ and $\mathtt{nei}_{\mathrm{C}}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$. Then, $\mathtt{AE}_{\mathrm{C}}(s)=\mathtt{nei}_{\mathrm{C}}(s)$.
@@
Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
@@
Let $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that $\mathrm{X}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Note that $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$ and $\mathrm{X}\neq\emptyset$. Then, $\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\neq\emptyset$, …, $\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\neq\emptyset$. Let $i$ be such that $1\leq i\leq n$ and $\sigma_{a_{i}}=\alpha_{x-u-\mathrm{X}_{i}}$. By the definition of $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})$, $\mathsf{out}_{a_{i}}(s,\sigma_{a_{i}})=\mathrm{X}_{i}$, $x=a_{i}$, and $u=s$. Then $\sigma_{a_{i}}=\alpha_{a_{i}-s-\mathrm{X}_{i}}$ and $\mathrm{X}_{i}\in\mathtt{nei}_{a_{i}}(s)$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$. Then, $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s)$.
@@
Let $\mathrm{X}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then, $\mathrm{X}=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}$ for some $\mathrm{X}_{1}\in\mathtt{nei}_{a_{1}}(s),\dots,\mathrm{X}_{n}\in\mathtt{nei}_{a_{n}}(s)$. Then, $\alpha_{a_{1}-s-\mathrm{X}_{1}}\dots,\alpha_{a_{n}-s-\mathrm{X}_{n}}\in\mathtt{AC}$. Note $\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})=\mathrm{X}_{1}$, $\dots$, $\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{n}$. Let $\sigma_{\mathrm{C}}=\alpha_{a_{1}-s-\mathrm{X}_{1}}\cup\dots\cup\alpha_{a_{n}-s-\mathrm{X}_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\alpha_{a_{1}-s-\mathrm{X}_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\alpha_{a_{n}-s-\mathrm{X}_{n}})=\mathrm{X}_{1}\cap\dots\cap\mathrm{X}_{n}=\mathrm{X}$. Note $\mathrm{X}$ is not empty. Then, $\mathrm{X}\in\mathtt{AE}_{\mathrm{C}}(s)$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model, and $\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of actual effectivity functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model. Assume $\mathtt{S}\textbf{-}\mathtt{AM}$ is $z$-representable by $\mathtt{S}\textbf{-}\mathtt{NM}$. Then, for all $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
@@ 5.1 Clear grand-coalition-first action models
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
We say that $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear if for all $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, if $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$, then $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$.
Intuitively, in clear grand-coalition-first action models, different action profiles have different outcome states. The following result indicates that in clear grand-coalition-first action models, different actions have different outcome states.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model, and $s\in\mathtt{ST}$.
@@
For any set $\mathrm{X}$ of states, we say that $\Delta\subseteq\mathcal{P}(\mathrm{X})$ is a general partition of $\mathrm{X}$ if (1) $\bigcup\Delta=\mathrm{X}$, and (2) for every $\mathrm{Y},\mathrm{Y}^{\prime}\in\Delta$, if $\mathrm{Y}\neq\mathrm{Y}^{\prime}$, then $\mathrm{Y}\cap\mathrm{Y}^{\prime}=\emptyset$. A general partition $\Delta$ is a partition if $\emptyset\notin\Delta$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model.
Then, if $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear, then for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is clear. Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. Note $\mathsf{out}_{\emptyset}(s,\emptyset)=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$. It is easy to show $\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Let $\mathrm{X}_{1},\mathrm{X}_{2}\in\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ such that $\mathrm{X}_{1}\neq\mathrm{X}_{2}$. Then, $\mathrm{X}_{1}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{1})$ for some $\sigma_{\mathrm{C}}^{1}\in\mathtt{JA}_{\mathrm{C}}$ and $\mathrm{X}_{2}=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}}^{2})$ for some $\sigma_{\mathrm{C}}^{2}\in\mathtt{JA}_{\mathrm{C}}$. Note $\sigma^{1}_{\mathrm{C}}\neq\sigma^{2}_{\mathrm{C}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})=\emptyset$, that is, $\mathrm{X}_{1}\cap\mathrm{X}_{2}=\emptyset$.
Note that the other direction of the statement of this fact might not hold. What follows is a counterexample.
Example 5.
Suppose $\mathtt{AG}=\{a\}$. Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a grand-coalition-first action model such that:
@@
It is easy to check for all $\mathrm{C}\subseteq\mathtt{AG}$, $\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$ is a general partition of $\mathsf{out}_{\emptyset}(s,\emptyset)$. However, $\mathtt{G}\textbf{-}\mathtt{AM}$ is not clear.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a clear grand-coalition-first action model.
@@
    for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.